OpenStudy (anonymous):

4 QUESTIONS FOR A MEDAL Simplify the radical expression √3 - 4 √3 a. 5√3 b. 5√6 c. -3√6 d. -3√3 Simplify the radical expression 2√6 + 3√96 a. 14√6 b.14√96 c.5√96 d. 50√6 Simplify the radical expression 6√5 - 8 √45 a. -2√50 b. -2√9 c. -48√225 d. -18√5 Simplify the expression√10 (√6 + 4) a.√6+4 b.√60+4√10 c.2√15 + 4√10 d.√16 + 4√10 pleaseeeee help meeeee fan & medal

OpenStudy (anonymous):

i guess nobody is going to help me :(

OpenStudy (owlcoffee):

I'll help you with the first one. if we have: \[\sqrt{3}-4\sqrt{3}\] if we take common factor "sqrt 3" we will have this: \[\sqrt{3}(1-4) \] we know that 1-4=-3 so then we'll replace: \[\sqrt{3}(-3)\] and since the multiplication is conmmutative: \[-3\sqrt{3}\]

OpenStudy (anonymous):

Ooooh okay thatnkssssss alot i didnt know how to break it down

OpenStudy (owlcoffee):

the main idea is simplifying them as much as possible.